4x^2+3y^2+z^2-8xy-6yz+4zx
1
Expert's answer
2021-09-05T18:38:39-0400

"4x^2+3y^2+z^2-8xy-6yz+4zx\\
=x^2+y^2+z^2-2xy-2yz+2zx+3x^2+2y^2-6xy-4yz\\
=(x-y+z)^2+2(x^2+y^2-2xy)+x^2-2xy-4yz\\
=(x-y+z)^2+2(x-y)^2+x^2-2xy-4yz"